OpenStudy (anonymous):

sqr root(3) +7 / 3 sqr root(3) -5

OpenStudy (anonymous):

2 questions ??

OpenStudy (anonymous):

\[\sqrt{3} +7/3\sqrt{3} -5\]

OpenStudy (anonymous):

no its 1

OpenStudy (anonymous):

ohk

OpenStudy (anonymous):

What's your question? Also is there any bracket?

OpenStudy (anonymous):

0.7735

OpenStudy (anonymous):

i got \[44+26\sqrt{3}\]

OpenStudy (anonymous):

\[22 + 13\sqrt{3}\]

OpenStudy (anonymous):

is the answer :)

OpenStudy (anonymous):